### v3

@@ 3.3 Phase I: Finding a feasible branch NLP
Regularization-based methods may not always find B-stationary points, but they can be efficient in finding (almost) feasible points. A Phase I method is proposed that combines solving LPECs with a regularization-based method Hoheisel2013 ; Kim2020 .
In principle, one can use any such method. In our exposition, we focus on Scholtes’ global relaxation method, which performs well in numerical experiments Kim2020 ; Hoheisel2013 ; Nurkanovic2024b . It solves a sequence of NLPs of the form:
@@
It is well-known that certifying infeasibility of a general nonconvex optimization problem is difficult. In many NLP solver implementations, if the globalization strategy cannot make sufficient progress toward a feasible point, a feasibility restoration phase is triggered, cf. (Fletcher2002, , Sec. 3.3.), (Waechter2006, , Sec 3.3.), and (Byrd2006, , Sec. 6). Thereby an auxiliary NLP – which minimizes the constraint violation – is solved Byrd2006 ; Fletcher2002 ; Waechter2006 . If its stationary point is infeasible for the original problem, the solver terminates. Infeasibility of the original problem is sufficient but not necessary for this outcome. This outcome is often called local infeasibility Waechter2006 , which is not a global certificate of infeasibility.
@@
with $a>1$. In the example, we set $a=1.1$. For larger values of $\tau$ both branches are feasible for the LPEC (15), but only the branch $x_{1}=0,x_{2}\geq 0$ is feasible for the MPEC. On the other hand, for smaller values of $\tau$, also in the LPEC, only the branch $d_{1}=0,\geq d_{2}\geq 0$ is feasible, thus the LPEC predicts only active set $({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$, which results in a feasible BNLP. Fig. 3 illustrates two sample solutions for $\mathrm{Reg}(\tau)$ and the corresponding $\mathrm{LPEC}(x^{*}(\tau),\rho)$. In this example, the smaller the parameter $a$ is the smaller $\tau$ must become.
(a) $\tau=1$.
(b) $\tau=1$.
@@
We show below in Corollary 4.1 that if the MPEC admits a feasible point, then for sufficiently small $\tau$, the solution of the LPEC corresponds to the index set of a feasible BNLP. Consequently, the while loop in Algorithm 2 has finite termination. In a practical implementation, a point is declared feasible if the constraint residual is below a certain tolerance tol_h. Formally, we require $h(x)\leq\texttt{tol}_{h}$, where $h(x)$ is the total infeasibility defined as:
@@
If $\mathrm{Reg}(\tau^{k})$ is infeasible (line 5), or its solution is feasible for the MPEC (line 7), the algorithm terminates. Otherwise, the solution $x^{*}(\tau^{k})$ is used to construct ${\mathrm{LPEC}}(x^{*}(\tau^{k}),\rho_{I})$. The fixed trust-region radius $\rho_{I}>0$ must be sufficiently large to ensure feasibility of this LPEC. Since a solution of $\mathrm{Reg}(\tau^{k})$ already satisfies the general inequality constraints, i.e., $c(x^{*}(\tau^{k}))\geq 0$, the smallest value of $\rho_{I}$ required for feasibility is $\rho_{I}>h_{\perp}(x^{*}(\tau^{k}))$.
The LPEC is solved only if this inequality holds (line 8). If it is not yet satisfied or the LPEC is infeasible, the homotopy parameter is reduced to obtain a better linearization point $x^{*}(\tau^{k})$ (line 15). If a solution to the LPEC is found, Algorithm 2 attempts to solve the corresponding $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ (line 11). If this is successful, a feasible point is obtained. Otherwise, the homotopy parameter is reduced to compute a better linearization point $x^{*}(\tau^{k+1})$.
@@ 3.4 Phase II: Computing a B-stationary point
4 MPECopt convergence proof
This section provides the convergence proof of the MPECopt method. Before stating the main results, we discuss the main assumption and some auxiliary results.
@@ 4.1 Assumptions
Most results in the previous section required only the MPEC-ACQ. For the convergence results in this section, we require a stronger assumption.
@@
    the classical MFCQ is satisfied at $x^{*}$ by the corresponding ${\mathrm{BNLP}}({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))$ for any partition $({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))\in\mathcal{P}(x^{*})$,
@@
    there exists a neighborhood $\mathcal{N}$ of $x^{*}$ and a constant $\bar{\tau}>0$ such that the classical MFCQ for Reg($\tau$) is satisfied at all $x\in\mathcal{N}\cap\Omega_{\mathrm{Reg}(\tau)}$ for all $\tau\leq\bar{\tau}$, where $\Omega_{\mathrm{Reg}(\tau)}$ is the feasible set of Reg($\tau$); and,
@@
    the tangent cones of the MPEC and the BNLPs satisfy:
@@
    Moreover, the left-hand side can be replaced by the MPEC-linearized feasible cone $\mathcal{F}_{\Omega}(x^{*})$, and the terms on the right-hand side by the classical linearized feasible cones of the BNLPs; cf. Eq. (40).
@@
Note that the globalization strategies of all listed NLP solvers enforce sufficient decrease conditions on every step to guarantee convergence to a stationary point. When starting from a feasible point, we assume that the step towards a stationary point satisfies a sufficient decrease conditions, which we summarize in Assumption 4.3 by inequality (27).
Even when starting from a feasible point, the NLP solver may fail to produce a point satisfying the globalization strategy (i.e., sufficiently feasible with improved objective). In such cases, a feasibility restoration phase is typically triggered Waechter2006 , possibly returning a point with higher objective. We did not observe this in any of our numerical experiments and therefore do not treat it further. A possible remedy is to add the constraint $f(x)\leq f(x^{k})+\epsilon$, with $\epsilon$ small, to the current BNLP.
@@ 4.2 Main convergence results
Proof. If $x$ is not B-stationary, then there exists at least one BNLP for which $x$ is not stationary. Denote by $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ the partition of ${\mathcal{I}}_{00}(x)$ associated with this BNLP.
By Proposition 4.1, the classical MFCQ holds for this BNLP. From this we draw two conclusions.
First, since $x$ is not B-stationary for this BNLP, Theorem 2.1 implies the existence of a direction $d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ such that $\nabla f(x)^{\top}d<0$. Moreover, because MFCQ holds, the tangent cone $\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ coincides with the classical linearized feasible cone of the BNLP, rendering (4) into an LP.
Second, considering the active sets for the BNLP at $x$, MFCQ implies the existence of a nonzero direction $s$ that points into the strict interior of the LP’s feasible set. In other words, the following set is nonempty:
@@
Combining these two facts yields the existence of $\epsilon>0$ and $s$ with $\|s\|_{\infty}=1$ such that (28) holds. ∎
Next, we show that an LPEC constructed at an almost feasible point can predict a feasible BNLP. In our implementation, such a point, and the corresponding LPEC are found in Algorithm 2.
@@
By requiring that the right-hand-side of  (31) and (32) are non-negative, we find that the linearized inequality constraints are feasible if $\rho$ satisfies:
@@
$\max\Big\{\frac{\delta}{\epsilon},\delta\Big\}\leq\rho\leq\min\Big\{\kappa_{2},\frac{\gamma_{1}}{\gamma_{2}}\Big\}$ (35)
the ${\mathrm{LPEC}}(\tilde{x},\rho)$ is feasible, since for a sufficiently small $\delta$ the interval (35) is nonempty.
@@
Remarkably, it follows from the last two results, that the LPEC in Phase I does not even need to be solved to optimality. It is sufficient to find a feasible point of the LPEC to identify a feasible BNLP.
Before we present the main convergence result, we prove that the inner loop in Algorithm 3 terminates in a finite number of iterations.
@@
Proof. If $d^{k,l}=0$ finite termination is trivial, so consider $d^{k,l}\neq 0$. Each step rejection in lines 10 or 12 multiplies $\rho$ by $\gamma^{\mathrm{L}}\in(0,1)$, hence after finitely many iterations there is an index $N\geq l$ with $\rho^{k,N}=(\gamma^{\mathrm{L}})^{N-l}\rho^{k,l}<\bar{\rho}$. For $\rho^{k,N}<\bar{\rho}$, due to Lemma 3.1 we can focus on the reduced LPEC (12). In this case, from the proof of Lemma 3.2 we have
$\mathcal{P}(x^{k})=\mathcal{P}(x^{k}+d^{k,N}),$
@@
    It computes a point $x^{*}$ that is B-stationary for the MPEC (1) in a finite number of iterations.
@@
Given a feasible point $x^{0}$, we show the outcomes of Algorithm 3.
If follows from Proposition 4.3 that the inner loop terminates in a finite number of iterations. Therefore, the outer loop it produces a sequence of points $\{x^{k}\}\subset\Omega$ with decreasing objective values $\{f^{k}\}$.
Consider some iterate $x^{k}$, which is a stationary point of some $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}),{\mathcal{I}}_{2}(x^{k}))$, and a $\rho^{k,l}>0$. If $d^{k,l}=0$ is a minimizer of $\mathrm{LPEC}(x^{k},\rho^{k,l})$, then a B-stationary point is found and the algorithm terminates in a finite number of iterations, which is outcome two.
We prove the occurrence of outcome three. Since $\{x^{k}\}\subset\Omega\subset X$ and $f$ is continuous while $X$ is compact, the sequence $\{f(x^{k})\}$ is bounded. By monotonicity, $\{f(x^{k})\}$ is convergent to some $f^{\infty}\in{\mathbb{R}}$. Consequently, $\{x^{k}\}$ has an accumulation point $x^{\infty}\in X$. Let $\{x^{k_{j}}\}$ be a subsequence converging to $x^{\infty}$. Suppose, for contradiction, that $x^{\infty}$ is not a B-stationary point. Then, by Assumption 4.3, there exists $\alpha^{k}>0$ such that $f(x^{k})-f(x^{k+1})\;\geq\;\alpha^{k}.$ Define $\bar{\alpha}:=\min_{k}\alpha^{k}>0$. It follows that for the subsequence, $f(x^{k_{j}})-f(x^{k_{j+1}})\;\geq\;\bar{\alpha},\quad\forall j\in\mathbb{N}.$ Summing over $j$ yields the telescoping bound
@@
We note that the assumption of isolated stationary points for finite termination is not overly restrictive. Under MPEC-MFCQ in Ass. 4.2, the classical MFCQ holds for the associated BNLPs. Moreover, if we additionally assume generalized second-order sufficient conditions at the BNLP stationary points, these points are isolated local minimizers (Forsgren2002, , Theorem 6.9).

### v4

@@ 3.3 Phase I: Finding a feasible branch NLP
Regularization-based methods may not always find B-stationary points, but they can be efficient in finding (almost) feasible points. We outline a Phase I method that combines solving LPECs with regularization-based methods. While any regularization-based method can be used in this framework, we focus in our exposition on Scholtes’ global relaxation method, which performs well in numerical experiments Kim2020 ; Hoheisel2013 ; Nurkanovic2024b . It solves a sequence of NLPs of the form:
@@
It is well-known that certifying infeasibility of a general nonconvex optimization problem is difficult. In many NLP solver implementations, if the globalization strategy cannot make sufficient progress toward a feasible point, a feasibility restoration phase is triggered, cf. (Fletcher2002, , Sec. 3.3.), (Waechter2006, , Sec 3.3.), and (Byrd2006, , Sec. 6). In this step, an auxiliary NLP that minimizes the constraint violation is solved Byrd2006 ; Fletcher2002 ; Waechter2006 . If its stationary point is infeasible for the original problem, the solver terminates. Infeasibility of the original problem is sufficient but not necessary for this outcome. This outcome is often called local infeasibility Waechter2006 , and it is not a global certificate of infeasibility.
@@
with $a>1$. In the example, we set $a=1.1$. For larger values of $\tau$ both branches are feasible for the LPEC (15), but only the branch $x_{1}=0,x_{2}\geq 0$ is feasible for the MPEC. On the other hand, for smaller values of $\tau$, also in the LPEC, only the branch $d_{1}=0,\ d_{2}\geq 0$ is feasible, thus the LPEC identifies only active sets $({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$, which results in a feasible BNLP. Fig. 3 illustrates two sample solutions for $\mathrm{Reg}(\tau)$ (left plots) and the corresponding $\mathrm{LPEC}(x^{*}(\tau),\rho)$ (right plots). In this example, the smaller the parameter $a$ is the smaller $\tau$ must become.
(a) $\tau=1.0$.
(b) $\tau=1.0$.
@@
We show below in Corollary 4.1 that if the MPEC admits a feasible point, then for sufficiently small $\tau$, then a feasible point of the LPEC yields a complementarity active set corresponding to a feasible BNLP. Consequently, the while loop in Algorithm 2 has finite termination. In a practical implementation, a point is declared feasible if the constraint residual is below a certain tolerance tolh. Formally, we require $h(x)\leq\texttt{tol}_{h}$, where $h(x)$ is the total infeasibility defined as:
@@
If $\mathrm{Reg}(\tau^{k})$ is infeasible (line 5), or its solution is feasible for the MPEC (line 7), the algorithm terminates. Otherwise, the solution $x^{*}(\tau^{k})$ is used to construct ${\mathrm{LPEC}}(x^{*}(\tau^{k}),\rho_{I})$. The fixed trust-region radius $\rho_{I}>0$ must be sufficiently large to ensure feasibility of this LPEC. Since a solution of $\mathrm{Reg}(\tau^{k})$ already satisfies the general inequality constraints, i.e., $c(x^{*}(\tau^{k}))\geq 0$, the smallest value of $\rho_{I}$ required for feasibility is $\rho_{I}>h_{\perp}(x^{*}(\tau^{k}))$. The LPEC is solved only if this inequality holds (line 8). If it is not yet satisfied or the LPEC is infeasible, the homotopy parameter is reduced to obtain a better linearization point $x^{*}(\tau^{k})$ (line 15). If a solution to the LPEC is found, Algorithm 2 attempts to solve the corresponding $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ (line 11). If this is successful, a feasible point is obtained. Otherwise, the homotopy parameter is reduced to compute a better linearization point $x^{*}(\tau^{k+1})$.
@@ 3.4 Phase II: Computing a B-stationary point
4 Convergence analysis
This section provides the convergence proof of the proposed method. Before stating the main results, we discuss the main assumption and some auxiliary results.
@@ 4.1 Assumptions
Some results in the previous section required only the MPEC-ACQ. For the convergence results in this section, we require a stronger assumption.
@@
    the usual MFCQ is satisfied at $x^{*}$ by the corresponding ${\mathrm{BNLP}}({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))$ for any partition $({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))\in\mathcal{P}(x^{*})$,
@@
    there exists a neighborhood $\mathcal{N}$ of $x^{*}$ and a constant $\bar{\tau}>0$ such that the usual MFCQ for Reg($\tau$) is satisfied at all $x\in\mathcal{N}\cap\Omega_{\mathrm{Reg}(\tau)}$ for all $\tau\leq\bar{\tau}$, where $\Omega_{\mathrm{Reg}(\tau)}$ is the feasible set of Reg($\tau$); and,
@@
    independently of Assumption 4.2, the tangent cones of the MPEC and the BNLPs satisfy
@@
    Moreover, under Assumption 4.2, the left-hand side can be replaced by the MPEC-linearized feasible cone $\mathcal{F}_{\Omega}(x^{*})$, and the terms on the right-hand side by the usual linearized feasible cones of the BNLPs; cf. Eq. (40).
@@
Note that the globalization strategies of all listed NLP solvers enforce sufficient decrease conditions on every step to guarantee convergence to a stationary point. When starting from a feasible point, we assume that the step towards a stationary point satisfies a sufficient decrease conditions, which we summarize in Assumption 4.3 by inequality (27). Even when starting from a feasible point, the NLP solver may fail to produce a point satisfying the globalization strategy (i.e., sufficiently feasible with improved objective). In such cases, a feasibility restoration phase is typically triggered Waechter2006 , possibly returning a point with higher objective. We did not observe this in any of our numerical experiments and therefore do not treat it further. A possible remedy is to add the constraint $f(x)\leq f(x^{k})+\epsilon$, with $\epsilon$ small, to the current BNLP.
@@ 4.2 Main convergence results
Proof. If $x$ is not B-stationary, then there exists at least one BNLP for which $x$ is not stationary. Denote by $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ the partition of ${\mathcal{I}}_{00}(x)$ associated with this BNLP. By Proposition 4.1, the usual MFCQ holds for this BNLP. From this we draw two conclusions.
First, since $x$ is not B-stationary for this BNLP, there exists a $d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ such that $\nabla f(x)^{\top}d<0$. Moreover, because MFCQ holds, the tangent cone $\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ coincides with the usual linearized feasible cone of the BNLP, rendering (4) into an LP.
Second, considering the active sets for the BNLP at $x$, MFCQ implies the existence of a nonzero direction $s$ that points into the strict interior of the LP’s (4) feasible set, i.e., the following set is nonempty:
@@
Next, let $\bar{s}\in S$. If $\nabla f(x)^{\top}\bar{s}<0$, set $s:=\bar{s}$. Otherwise, pick any $d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ with $\nabla f(x)^{\top}d<0$. Because $\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ is a cone then for every $\lambda>0$ we can define a new descent direction $d_{\lambda}:=\lambda d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$. Next, define the vector $s(t):=\bar{s}+td_{\lambda}$. Since $\bar{s}\in S$, there exists $t_{0}>0$ such that $s(t)\in S$ for all $t\in[0,t_{0}]$. Moreover,
$\nabla f(x)^{\top}s(t)=\nabla f(x)^{\top}\bar{s}+t\,\nabla f(x)^{\top}d_{\lambda}=\nabla f(x)^{\top}\bar{s}+t\lambda\,\nabla f(x)^{\top}d.$
Choose $\lambda$ so large that $0<\frac{\nabla f(x)^{\top}\bar{s}}{-\lambda\nabla f(x)^{\top}d}<t_{0}$, and then take any $t\in\big(\frac{\nabla f(x)^{\top}\bar{s}}{-\lambda\nabla f(x)^{\top}d},\,t_{0}\big]$. This yields $s(t)\in S$ and $\nabla f(x)^{\top}s(t)<0$. Fix such $(\lambda,t)$ and set $s:=s(t)$.
Because inequalities are strict for $s(t)$, define $\epsilon_{1}:=-\nabla f(x)^{\top}s,\ \epsilon_{2}:=\min_{i\in{\mathcal{A}}(x)}\nabla c_{i}(x)^{\top}s,\ \epsilon_{3}:=\min(\min_{i\in{\mathcal{D}}_{1}(x)}s_{2,i},\!\min_{i\in{\mathcal{D}}_{2}(x)}s_{1,i}),$ which are all positive, and set $\bar{\epsilon}:=\min\{\epsilon_{1},\epsilon_{2},\epsilon_{3}\}>0$. With $\hat{s}:=s/\|s\|_{\infty}$ and $\epsilon:=\bar{\epsilon}/\|s\|_{\infty}$, the vector $\hat{s}$ satisfies (28).
Next, we show that an LPEC constructed at an almost feasible point can “predict” a feasible BNLP. In our implementation, such a point, and the corresponding LPEC are found in Algorithm 2.
@@
By requiring that the right-hand-side of (31) and (32) are nonnegative, we find that the linearized inequality constraints are feasible if $\rho$ satisfies:
@@
$\max\Big\{\frac{\delta}{\epsilon},\delta\Big\}\leq\rho\leq\min\Big\{\kappa_{2},\frac{\gamma_{1}}{\gamma_{2}}\Big\}.$ (35)
Since there always exists a $\varepsilon>0$ for every $x^{*}\in\Omega$ (see proof of Proposition 4.2), the term $\lim\limits_{\delta\to 0}\frac{\varepsilon}{\delta}=0$ is well-defined. Therefore, the ${\mathrm{LPEC}}(\tilde{x},\rho)$ is feasible, since for a sufficiently small $\delta$ the interval (35) is nonempty.
@@
Remarkably, it follows from the last two results, that the LPEC in Phase I does not even need to be solved to optimality. It is sufficient to find a feasible point of the LPEC to identify a feasible BNLP. Before we present the main convergence result, we prove that the inner loop in Algorithm 3 terminates in a finite number of iterations.
@@
Proof. If $d^{k,l}=0$ finite termination is trivial, so consider $d^{k,l}\neq 0$. Each step rejection in lines 10 or 12 multiplies $\rho$ by $\gamma^{\mathrm{L}}\in(0,1)$, hence after finitely many iterations there is an index $N\geq l$ with $\rho^{k,N}=(\gamma^{\mathrm{L}})^{N-l}\rho^{k,l}<\bar{\rho}$. For $\rho^{k,N}<\bar{\rho}$, due to Lemma 3.1 we can focus on the reduced LPEC (12). In this case, because $\|d^{k,N}\|_{\infty}<\bar{\rho}$ is sufficiently small, from the proof of Lemma 3.2 we have
$\mathcal{P}(x^{k}+d^{k,N})\subseteq\mathcal{P}(x^{k}),$
@@
    It computes a point $x^{*}$ that is B-stationary for the MPEC (1) after solving finitely many BNLPs and LPECs.
@@
Given a feasible point $x^{0}$, we discuss the outcomes of Algorithm 3. If follows from Proposition 4.3 that the inner loop terminates in a finite number of iterations. Therefore, the outer loop it produces a sequence of points $\{x^{k}\}\subset\Omega$ with decreasing objective values $\{f^{k}\}$. Consider some iterate $x^{k}$, which is a stationary point of some $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}),{\mathcal{I}}_{2}(x^{k}))$, and a $\rho^{k,l}>0$. If $d^{k,l}=0$ is a minimizer of $\mathrm{LPEC}(x^{k},\rho^{k,l})$, then a B-stationary point is found and the algorithm terminates in a finite number of iterations, which is outcome two.
We prove the occurrence of outcome three. Since $\{x^{k}\}\subset\Omega\subset X$ and $f$ is continuous while $X$ is compact, the sequence $\{f(x^{k})\}$ is bounded. By monotonicity, $\{f(x^{k})\}$ is convergent to some $f^{\infty}\in{\mathbb{R}}$. Consequently, $\{x^{k}\}$ has an accumulation point $x^{\infty}\in X$. Let $\{x^{k_{j}}\}$ be a subsequence converging to $x^{\infty}$. Suppose, for contradiction, that $x^{\infty}$ is not a B-stationary point. Recall that each $x^{k},k>0$ is a stationary point of a BNLP, and by Assumption 4.3, for each $k\geq 0$ there exist an $\alpha^{k}>0$ such that $f(x^{k})-f(x^{k+1})\geq\alpha^{k}>0.$ Therefore, we can define an uniform bound $\bar{\alpha}:=\min_{k}\alpha^{k}>0$. It follows that for the subsequence, $f(x^{k_{j}})-f(x^{k_{j+1}})\;\geq\;\bar{\alpha},\quad\forall j\in\mathbb{N}.$ Summing over $j$ yields the telescoping bound
@@
We note that the assumption of isolated stationary points for finite termination is not overly restrictive. Under MPEC-MFCQ in Ass. 4.2, the usual MFCQ holds for the associated BNLPs. Moreover, if we additionally assume generalized second-order sufficient conditions at the BNLP stationary points, these points are isolated local minimizers (Forsgren2002, , Theorem 6.9).